# A Survey of World Models for Autonomous Driving

### v3

@@ 3.2.1 Learning-based Planning
TABLE IV: Overview of representative rule-based motion-planning algorithms (§3.2.2).
@@
Car-following Models IDM [154] Phys. Rev. E’00 Closed-form ordinary-differential law Gap + Speed + Relative speed Longitudinal acceleration
IDM Roll-out Sampling [155] CoRL’23 Batched IDM roll-outs with cost ranking Gap + Speed + Relative speed Minimum-cost acceleration
@@
Sampling-based Random RRT [156] Tech. Rep.’98 Tree search with random exploration State + Map + Obstacles Feasible path
RRT-Connect [157] ICRA’00 Bi-directional RRT connection State + Map + Obstacles Feasible path
RRT/RRG [158, 159] RSS’10/IJRR’11 Rewiring optimal RRT for asymptotic optimality State + Map + Obstacles Near-optimal path
[…]
Conformal Lattice [161] ICRA’11 Highway-aligned lattice search Ego + State + Map Trajectory
Discrete Terminal Manifold [162] IJRR’12 Polynomial library with A* search Ego + State + Map Time-critical trajectory
@@
Continuous Optimisation Apollo EM Planner [163] arXiv’18 Alternating quadratic programming + dynamic programming Frenet path + Road corridor Smooth trajectory
Continuous-Curvature Spline [164] IV’14 Bézier-spline quadratic programming with jerk minimisation Obstacle corridor Smooth trajectory
@@
Artificial Potential-Field Classic APF [165, 166] CISM’78/IJRR’86 Attractive–repulsive potential field Goal pose + Obstacles Steering force
@@
Newton-PF [167] T-RO’06 Potential-field with Newton-based update Goal pose + Obstacles Steering force
@@
Hybrid Sigmoid-PF [168, 169] Sensors’10/IV’17 Potential field with sigmoid blending Goal pose + Obstacles Steering command
@@
Torque-PD PF [170] IV’15 Potential field within proportional–derivative steering Potential-field value Steering torque
Summary. Table III shows compares major learning-based planning paradigms. Several observations can be drawn:
[…]
3.2.2 Rule-based Planning
Rule-based planners endure because they guarantee safety and are easy to audit [171, 172]. They treat motion as a deterministic optimisation over explicit dynamics and hand-crafted costs, yielding transparent behaviour. Four core families dominate: car-following laws (e.g., IDM), sampling planners (RRT, state lattice), model-predictive control, and potential-field navigation. Modern stacks combine these blocks for interpretability and safety, yet modular fusion can lose global optimality and fixed heuristics struggle in dense, fast-changing traffic.
Car-following Models. Physics-inspired longitudinal models translate the instantaneous gap, speed and closing rate into throttle or brake commands. The Intelligent Driver Model and its calibrated siblings capture free-flow, synchronized and stop‐and-go regimes with a single continuous formula [154]. Recent work samples multiple IDM roll-outs and selects the minimum-cost, safety-compliant candidate to improve comfort without losing closed-form guarantees [155]. Because these laws are purely algebraic, production-grade Adaptive Cruise Control (ACC) systems can execute them at tens-to-hundreds Hz frequencies. However, their inherently short prediction horizon is a key limitation, so modern planners augment analytic commands with map priors or predictive MPC to anticipate intersections and cut-in maneuvers better.
Sampling-based Planners. Sampling-based planners probe the vehicle’s configuration space with a sequence of samples, accept only collision-free continuations, and then rank feasible motion sequences with handcrafted costs. They split into two sub-families that trade optimality for speed in different ways. Random sampling methods grow a tree or roadmap toward stochastically chosen states, guaranteeing probabilistic completeness and – when augmented with rewiring – asymptotic optimality. Classic Rapidly-Exploring Random Trees (RRT) [156, 157] realise millisecond-level feasibility checks for single-query problems. Their optimal variants, RRT and RRG, refine paths online by incremental rewiring [158, 159]. Closed-loop extensions embed vehicle dynamics for high-speed manoeuvres, as demonstrated on MIT’s “Talos” car in the DARPA Urban Challenge [173, 174] and in multi-agent settings [175]. Deterministic sampling methods pre-tile the state–time space with kinematically feasible motion primitives, then apply graph search to pick the lowest-cost, collision-free sequence. Spatiotemporal lattices and their conformal highway variant capture curvature and speed limits explicitly [160, 161]. Polynomial libraries (cubic, quintic, B-spline) further reduce run-time by storing analytically smooth segments that satisfy dynamic bounds [176, 177]. Discretised terminal manifolds shrink search depth for time-critical street scenarios [162], and the Stanford “Junior” entry shows how such libraries integrate into a full urban-driving stack [178]. Because the candidate set is finite, deterministic planners sacrifice global optimality for predictable timing, making them popular for production autonomous vehicle stacks.
Continuous Optimisation. Continuous optimisation planners rebuild vehicle paths in continuous space each control cycle, casting curvature and speed as variables in a quadratic or nonlinear program constrained by dynamics and obstacle corridors. Warm-starting with the previous solution yields smooth, near-optimal trajectories without discrete motion libraries. Apollo’s Expectation–Maximisation planner alternates lateral smoothing with longitudinal speed optimisation, whereas Rastelli’s continuous-curvature scheme minimises jerk and curvature variation along Bézier splines. Both show that convex or near-convex formulations can meet real-time demands while overcoming the sub-optimality and roughness of sampling-based schemes, though at the cost of continual optimisation [163, 164].
Artificial Potential-field Navigation. Attractive and repulsive potentials characterise the goal as an energy well and obstacles as peaks, letting a robot follow the negative gradient in real time [165, 166]. Newton-style updates hasten escape from saddle points [167], and adaptive gain tuning lessens oscillations on public roads [169]. Blending the field with sigmoid curves limits curvature spikes in dense traffic [168], while embedding the gradient in a torque-based proportional–derivative steering loop unifies motion planning with low-level control [170]. Although local minima persist, coupling potential-field guidance with higher-level sampling or MPC preserves rapid collision avoidance without forfeiting global optimality [171, 172].
Summary. Table IV compares four major rule-based motion-planning paradigms. Several observations follow:
[…]
Figure 3: The evolution of interaction3.3). (a) Open-loop regime synthesises scenes from logged data under static conditions; the generator memorises scenarios but never reacts to new actions. (b) Uncontrollable closed-loop regime unifies perception, prediction and planning in an auto-regressive loop, yet latent physics remain fixed, so users cannot inject rules or rare events. (c) Controllable closed-loop regime adds editable 4D worlds, occupancy control and feedback/backward signals, enabling a drive agent to interact safely with a fully testable simulator.
3.2.3 Search-based Planning
@@
Search-based planners discretise the vehicle’s high-dimensional, continuous state-space into a graph whose edges are motion primitives – short, pre-computed, kinematically feasible maneuvers. Classical graph-search techniques such as Dijkstra [179, 180] and the A* family [181, 182] then find the minimum-cost path according to heuristics that balance safety, smoothness, and progress. Vehicle constraints are embedded in the primitive library; Hybrid-State A* extends this idea by blending continuous headings with discrete cells to satisfy non-holonomic bicycle-model dynamics [183].
@@
The approach offers strong optimality guarantees and explicit handling of traffic rules and obstacles, but suffers from the resolution-complexity trade-off: finer grids or larger graphs improve path quality and continuity yet raise computational load exponentially, jeopardising real-time performance. Recent work tackles this bottleneck with (i) model-predictive A* that employs relaxed cost-to-go maps for long-horizon searches in milliseconds [184], (ii) adaptive fidelity models that switch between simple longitudinal or drift dynamics to match task urgency [185], (iii) multi-heuristic search that fuses several admissible and inadmissible heuristics to accelerate exploration in cluttered layouts [186], and (iv) Task-and-Motion Planning hybrids that sample primitives from multiple local models while maintaining optimality guarantees in nonlinear, unstable regimes [187]. Together these advances extend search-based planning from structured highway settings to complex urban and off-road scenarios while retaining its interpretable, graph-optimal foundation.
3.3 Interaction between Planning and Prediction
@@
Generative Methods and Open-loop Regime. Recent generative approaches (e.g., DriveGAN [188], MagicDrive [189], and DriveDreamer [10]) narrow the sim-to-real gap by focusing on photorealism and synthesizing editable street-scene videos that substantially enrich training corpora; nevertheless, they merely replay pre-sampled futures without responding to online control inputs, thus confining research to a data-centric, open-loop paradigm that breaks the causal link between present actions and subsequent observations. TrafficGen [190] and LCTGen [191] generate traffic scenes from vector logs and natural-language prompts, respectively, while RealGen [192] retrieves and recombines template behaviours to create tag-specific or crash-critical scenarios. All three remain strictly open-loop, replaying fixed futures that ignore online control inputs. NAVSIM [193] shares this non-reactive paradigm, seeking only to refine open-loop metrics; it cannot inject rare events or alter traffic rules, offering merely quasi-closed-loop scores. These limitations draw growing criticism, as open-loop benchmarks can overstate safety and performance when an agent’s actions never influence the futures being evaluated.
@@
${}_{\!}$ ${}_{\!}$ ${}_{\!}$ ${}_{\!}$ ${}_{\!}$ ${}_{\!}$ Regime. Auto-regressive driving world models [15, 46, 84, 194, 195, 196, 88, 85, 89, 197, 198] have quickly progressed from trajectory-only roll-outs to multi-sensor, long-horizon simulators that generate futures conditioned on ego actions, unifying prediction and planning in a single generative loop. This paradigm supplies cheap, diverse data, exposes planners to their own downstream consequences, and captures uncertainty more richly than fixed-log, open-loop corpora [15, 46, 199]. However, the latent physics are opaque: users cannot freely edit traffic rules, inject rare events, adjust replay speed, or verify safety guarantees, and compounding distribution drift still threatens reliability, leaving current systems in an uncontrollable closed-loop that hampers rigorous safety-critical evaluation [21].
@@
Prior auto-regressive models relied primarily on image [15, 46, 84]. Recent work [85, 89] has integrated 3D occupancy grids into environmental modeling, providing a more detailed and robust representation of spatial dynamics. These grids enable accurate handling of occlusions and spatial relationships in dense urban scenes. However, their use in long-sequence video generation remains limited by the computational complexity of processing and maintaining large-scale 3D data over extended periods.
@@
Controllable Closed-loop Regime. Research has progressed from script-driven traffic engines, through game-engine simulators with sensor realism, to neural and hybrid platforms that merge photoreal rendering with fully editable actor control. Early microscopic engines such as SUMO [200] exposed per-vehicle APIs for spawning, deletion and IDM/MOBIL/RL control, enabling fine-grained, reproducible experiments; Waymax [201] later accelerated log replay with JAX for rapid, sensor-free policy iteration, yet neither rendered raw sensor data. To fill this gap, CARLA [49] grafted Unreal-Engine visuals and configurable sensors onto a physics core, while MetaDrive [202] traded photoreal texture for procedural roads and kHz physics to support fast domain randomisation. Both deliver closed-loop images but still cover limited geographies and fall short of real-world visual fidelity, restricting direct transfer to models trained on natural data. Neural-hybrid simulators then appeared. UniSim [203] learns photoreal feature grids from a single drive, and OASim [204] re-poses vehicles with lighting-consistent implicit rendering, yet both remain bounded by their source clips. Fidelity and interactivity keeps expanding: Sky-Drive [205] adds socially aware multi-agent interactions; DriveArena [29] converts 2D traffic sketches to photoreal video; DrivingSphere [206] upgrades to 4D occupancy for geometrically accurate stress tests. The LimSim family scales closed-loop control to city networks. LimSim [207] offers distributed, scriptable traffic; LimSim++ [208] adds road-rule scripts and vision hooks; and the latest LIMSIM Series [209] provides modular APIs for covering map topology, human-style bots and area-of-interest acceleration, allowing users to inject hazards, swap policies and tune metrics for rigorous, interactive safety validation.
@@
4 Data and Training Paradigms
@@
4.1 Self-Supervised Learning for World Models
Self-supervised world-model research for autonomous driving uses 2D images or raw LiDAR scans to automatically generate supervision, cutting 3D annotation costs. Rapid progress is now reshaping the field around 3D occupancy grids. Because dense voxel annotation is laborious, many groups pivoted to cheaper 2D supervision. RenderOcc [210], SelfOcc [211], OccNeRF [212] and related NeRF-style systems lift multi-view images into volumetric space, back-propagating depth or semantic losses so the network “hallucinates” occupancy; these image-only routes save annotation cost but lag LiDAR-centric baselines in fidelity and fine geometry. A complementary branch exploits the LiDAR stream itself: UnO [31] and EO [117] derive self-labels by contrasting predicted 4D fields with future scans, while RenderWorld [91], UniPAD [213], and PreWorld [82] couple differentiable rendering with multi-view cameras to fuse 2D texture and 3D structure; ViDAR [20] further aligns video patches to point clouds for stronger temporal motion priors. Tokenisation and diffusion bring a third flavour: COPILOT4D [214] and BEVWorld [32] discretise multimodal data into vocabulary codes, then learn discrete diffusion to regenerate future scenes, capturing complex distributions without any human annotation, and AD-L-JEPA [215] eliminates both contrastive and generative heads, directly forecasting bird’s-eye embeddings for faster pre-training.

### v4

@@ 3.2.2. Rule-based Planning
Table 4. Overview of representative rule-based motion-planning algorithms (§3.2.2).
@@
Car-following Models IDM (Treiber et al., 2000) Phys. Rev. E’00 Closed-form ordinary-differential law Gap + Speed + Relative speed Longitudinal acceleration
IDM Roll-out Sampling (Dauner et al., 2023) CoRL’23 Batched IDM roll-outs + cost ranking Gap + Speed + Relative speed Acceleration + Rollout
@@
Sampling-based Random RRT (LaValle, 1998) Tech. Rep.’98 Tree search with random exploration State + Map + Obstacles Feasible path
RRT-Connect (Kuffner and LaValle, 2000) ICRA’$0$ Bi-directional RRT connection State + Map + Obstacles Feasible path
RRT/RRG (Karaman and Frazzoli, 2010, 2011) RSS’10/IJRR’11 Rewiring optimal RRT for asymptotic optimality State + Map + Obstacles Near-optimal path
[1pt/1pt] Sampling-based Deterministic Spatiotemporal Lattice (Ziegler and Stiller, 2009) IROS’09 Pre-tiled spatiotemporal lattice search Ego + State + Map Trajectory
Conformal Lattice (McNaughton et al., 2011) ICRA’11 Highway-aligned lattice search Ego + State + Map Trajectory
Discrete Terminal Manifold (Werling et al., 2012) IJRR’12 Polynomial trajectory library + Cost ranking Ego + State + Map Time-critical trajectory
@@
Continuous Optimisation Apollo EM Planner (Fan et al., 2018) arXiv’18 Alternating DP + spline-QP Frenet reference line + Feasible tunnel Smooth trajectory
Continuous-Curvature Spline (Pérez Rastelli et al., 2014) IV’14 Bézier-spline continuous-curvature planning + Jerk-aware cost Ego + Map Smooth trajectory
@@
Artificial Potential-Field Classic APF (Khatib and Le Maitre, 1978; Khatib, 1986) CISM’78/IJRR’86 Attractive-repulsive potential field Goal pose + Obstacles Steering command
@@
Newton-PF (Ren et al., 2006) T-RO’06 Potential-field with Newton-based update Goal pose + Obstacles Steering command
@@
Hybrid Sigmoid-PF (Lu et al., 2020; Bounini et al., 2017) Sensors’20/IV’17 Potential field with sigmoid blending Goal pose + Obstacles Steering command
@@
Torque-PD PF (Galceran et al., 2015) IV’15 Potential field within proportional–derivative steering Potential-field value Steering torque
3.2.3. Search-based Planning
@@
Search-based planners discretise the vehicle’s high-dimensional, continuous state-space into a graph whose edges are motion primitives – short, pre-computed, kinematically feasible maneuvers. Classical graph-search techniques such as Dijkstra (Gu et al., 2016; Bohren et al., 2009) and the A* family (Kammel et al., 2009; Boroujeni et al., 2017) then find the minimum-cost path according to heuristics that balance safety, smoothness, and progress. Vehicle constraints are embedded in the primitive library; Hybrid-State A* extends this idea by blending continuous headings with discrete cells to satisfy non-holonomic bicycle-model dynamics (Dolgov et al., 2008).
@@
The approach offers strong optimality guarantees and explicit handling of traffic rules and obstacles, but suffers from the resolution-complexity trade-off: finer grids or larger graphs improve path quality and continuity yet raise computational load exponentially, jeopardising real-time performance. Recent work tackles this bottleneck with (i) model-predictive A* that employs relaxed cost-to-go maps for long-horizon searches in milliseconds (Ajanovic et al., 2018), (ii) adaptive fidelity models that switch between simple longitudinal or drift dynamics to match task urgency (Ajanovic, 2019), (iii) multi-heuristic search that fuses several admissible and inadmissible heuristics to accelerate exploration in cluttered layouts (Adabala and Ajanović, 2023), and (iv) Task-and-Motion Planning hybrids that sample primitives from multiple local models while maintaining optimality guarantees in nonlinear, unstable regimes (Ajanović et al., 2023). Together these advances extend search-based planning from structured highway settings to complex urban and off-road scenarios while retaining its interpretable, graph-optimal foundation.
3.3. Interaction between Planning and Prediction
@@
Generative Methods and Open-loop Regime. Recent generative approaches (e.g., DriveGAN (Kim et al., 2021), MagicDrive (Gao et al., 2024b), and DriveDreamer (Wang et al., 2024c)) narrow the sim-to-real gap by focusing on photorealism and synthesizing editable street-scene videos that substantially enrich training corpora; nevertheless, they merely replay pre-sampled futures without responding to online control inputs, thus confining research to a data-centric, open-loop paradigm that breaks the causal link between present actions and subsequent observations. TrafficGen (Feng et al., 2023a) and LCTGen (Tan et al., 2023) generate traffic scenes from vector logs and natural-language prompts, respectively, while RealGen (Ding et al., 2024a) retrieves and recombines template behaviours to create tag-specific or crash-critical scenarios. All three remain strictly open-loop, replaying fixed futures that ignore online control inputs. NAVSIM (Dauner et al., 2024) shares this non-reactive paradigm, seeking only to refine open-loop metrics; it cannot inject rare events or alter traffic rules, offering merely quasi-closed-loop scores. These limitations draw growing criticism, as open-loop benchmarks can overstate safety and performance when an agent’s actions never influence the futures being evaluated.
Figure 3. The evolution of interaction3.3). (a) Open-loop regime synthesises scenes from logged data under static conditions; the generator memorises scenarios but never reacts to new actions. (b) Uncontrollable closed-loop regime unifies perception, prediction and planning in an auto-regressive loop, yet latent physics remain fixed, so users cannot inject rules or rare events. (c) Controllable closed-loop regime adds editable 4D worlds, occupancy control and feedback/backward signals, enabling a drive agent to interact safely with a fully testable simulator.
@@
Auto-regressive World Models and Uncontrollable Closed-loop Regime. Auto-regressive driving world models (Hu et al., 2023a; Gao et al., 2024c; Bogdoll et al., 2023; Zhao et al., 2024; Wu et al., 2024a; Yang et al., 2025a; Zheng et al., 2024; Yang et al., 2025c; Zheng et al., 2025a; Wei et al., 2024; Suo et al., 2021; Zhang et al., 2023a, 2025; Feng et al., 2025) have quickly progressed from trajectory-only roll-outs to multi-sensor, long-horizon simulators that generate futures conditioned on ego actions, unifying prediction and planning in a single generative loop. This paradigm supplies cheap, diverse data, exposes planners to their own downstream consequences, and captures uncertainty more richly than fixed-log, open-loop corpora (Hu et al., 2023a; Gao et al., 2024c; Chen et al., 2025a). However, the latent physics are opaque: users cannot freely edit traffic rules, inject rare events, adjust replay speed, or verify safety guarantees, and compounding distribution drift still threatens reliability, leaving current systems in an uncontrollable closed-loop that hampers rigorous safety-critical evaluation (Min et al., 2024b).
@@
Some autoregressive models rely primarily on images (Hu et al., 2023a; Gao et al., 2024c; Bogdoll et al., 2023; Mousakhan et al., 2025). Another line of autoregressive models (Zheng et al., 2025a; Wei et al., 2024) has integrated 3D occupancy grids into environmental modeling, providing a more detailed and robust representation of spatial dynamics. These grids enable accurate handling of occlusions and spatial relationships in dense urban scenes. However, their use in long-sequence video generation remains limited by the computational complexity of processing and maintaining large-scale 3D data over extended periods. Concurrently, state-space diffusion at the agent/traffic-light level scales to trip-level city simulation(Tan et al., 2025), complementing image-/occupancy-based rollouts with controllable mid-level semantics.
@@
Controllable Closed-loop Regime. Research has progressed from script-driven traffic engines, through game-engine simulators with sensor realism, to neural and hybrid platforms that merge photoreal rendering with fully editable actor control. Early microscopic engines such as SUMO (Krajzewicz et al., 2012) exposed per-vehicle APIs for spawning, deletion and IDM/MOBIL/RL control, enabling fine-grained, reproducible experiments; Waymax (Gulino et al., 2024) later accelerated log replay with JAX for rapid, sensor-free policy iteration, yet neither rendered raw sensor data. To fill this gap, CARLA (Dosovitskiy et al., 2017b) grafted Unreal-Engine visuals and configurable sensors onto a physics core, while MetaDrive (Li et al., 2022b) traded photoreal texture for procedural roads and kHz physics to support fast domain randomisation. Both deliver closed-loop images but still cover limited geographies and fall short of real-world visual fidelity, restricting direct transfer to models trained on natural data. Neural-hybrid simulators then appeared. UniSim (Yang et al., 2023a) learns photoreal feature grids from a single drive, and OASim (Yan et al., 2024) re-poses vehicles with lighting-consistent implicit rendering, yet both remain bounded by their source clips. Dreamland (Mo et al., 2025) links a physics simulator with a large video generator via a layered world abstraction, mapping simulator states to editable photoreal imagery and boosting controllability and realism for configurable closed-loop world creation. Fidelity and interactivity keeps expanding: Sky-Drive (Huang et al., 2025a) adds socially aware multi-agent interactions; DriveArena (Yang et al., 2024d) converts 2D traffic sketches to photoreal video; DrivingSphere (Yan et al., 2025b) upgrades to 4D occupancy for geometrically accurate stress tests. The LimSim family scales closed-loop control to city networks. LimSim (Wenl et al., 2023) offers distributed, scriptable traffic; LimSim++ (Fu et al., 2024a) adds road-rule scripts and vision hooks; and the latest LIMSIM Series (Fu et al., 2025) provides modular APIs for covering map topology, human-style bots and area-of-interest acceleration, allowing users to inject hazards, swap policies and tune metrics for rigorous, interactive safety validation.
Beyond simulator-centric loops, recent works realize controllable closed loops inside the latent space. World4Drive (Zheng et al., 2025c) builds an intention-aware latent world model, rolls out intent-conditioned futures and uses a world-model selector to choose the best trajector. LAW (Li et al., 2025a) learns a plan-conditioned latent world model; future-feature prediction supervises reps and waypoint planning.
@@
4. Data and Training Paradigms
@@
Figure 4. (a) Self-supervised world models lift multi-view images into 3D volumes, forecast future grids, and learn from 2D renders plus occupancy cues, reducing labels (§4.1). (b) Large-scale pre-training on visual point-cloud sequences yields a single backbone that adapts to detection, tracking, mapping, occupancy, and planning (§4.2).
4.1. Self-Supervised Learning for World Models
Self-supervised world-model research for autonomous driving uses 2D images (Jin et al., 2025) or raw LiDAR scans to automatically generate supervision, cutting 3D annotation costs. Rapid progress is now reshaping the field around 3D occupancy grids. Because dense voxel annotation is laborious, many groups pivoted to cheaper 2D supervision. RenderOcc (Pan et al., 2024), SelfOcc (Huang et al., 2024b), OccNeRF (Zhang et al., 2023c), H3O (Shi et al., 2025a), World4Drive (Zheng et al., 2025c), and related NeRF-style systems lift multi-view images into volumetric space, back-propagating depth or semantic losses; these image-only routes save annotation cost but lag LiDAR-centric baselines in fidelity and fine geometry. A complementary branch exploits the LiDAR stream itself: UnO (Agro et al., 2024) and EO (Khurana et al., 2022) derive self-labels by contrasting predicted 4D fields with future scans, while RenderWorld (Yan et al., 2025a), UniPAD (Yang et al., 2024e), and PreWorld (Li et al., 2025b) couple differentiable rendering with multi-view cameras to fuse 2D texture and 3D structure; ViDAR (Yang et al., 2024a) further aligns video patches to point clouds for stronger temporal motion priors. Tokenisation and diffusion bring a third flavour: COPILOT4D (Zhang et al., 2023b) and BEVWorld (Zhang et al., 2024a) discretise multimodal data into vocabulary codes, then learn discrete diffusion to regenerate future scenes, capturing complex distributions without any human annotation, and AD-L-JEPA (Zhu et al., 2025a) eliminates both contrastive and generative heads, directly forecasting bird’s-eye embeddings for faster pre-training.